OpenStudy (anonymous):

Which equation has a solution of -2?

Parth (parthkohli):

\[x + 2 = 0 \\ x + 3 = 1\\x+4=2\\\vdots\\2x=-4\]Infinite equations like these exist.

Parth (parthkohli):

But you gotta be specific, okay?

OpenStudy (anonymous):

ok